before signing a lease, a company reports total assets of $500,000 and total liability $300,000. The company then sign a 30-month lease for equipment with payment of $922.21 each month. The lease payment have a present value of $25,000. After recording the inception of the lease, The company would report which of the following?

a. asset of 527,666.30 and total libiablity of 325,000.00

b. asset of 525,000.20 and total libiablity of 327,666.30

c. asset of 527,666.30 and total libiablity of 327,666.30

d. asset of 525,000.00 and total libiablity of 325,000.00
Answer #1

Assets = 500000+25000 = $525000

Liabilities = 300000 + 25000 = $325000

Option d. is correct answer.